40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 7 20:25:14 2016, Koji, Update, CDS, Power-cycled c1psl and c1iool0 
    Reply  Mon Oct 10 11:48:05 2016, gautam, Update, CDS, Power-cycled c1susaux, realigned PMC, spots centered on WFS1 and WFS2 
Message ID: 12542     Entry time: Mon Oct 10 11:48:05 2016     In reply to: 12539
Author: gautam 
Type: Update 
Category: CDS 
Subject: Power-cycled c1susaux, realigned PMC, spots centered on WFS1 and WFS2 

[Koji, Gautam]

We did the following today morning:

  1. I re-aligned the PMC - transmission level on the scope on the PSL table is now ~0.72V which is around what I remember it being
  2. The spot had fallen off WFS 2 - so we froze the output of the MC WFS servo, and turned the servo off. Then we went to the table to re-center the spot on the WFS. The alignment had drifted quite a bit on WFS2, and so we had to change the scale on the grid on the MEDM screen to +/-10 (from +/- 1) to find the spot and re-center it using the steering mirror immediately before the WFS. It would appear that the dark offsets are different on WFS1 and WFS2, so the "SUM" reads ~2.5 on WFS1 and ~0.3 on WFS2 when the spots are well centered
  3. Coming back to the control room, we ran the WFSoffsets script and turned on the WFS servo again. Trying to run the relief servo, we were confronted by an error message that c1susaux needed to be power cycled (again). This is of course the slow machine that the ITMX suspension is controlled by, and in the past, power cycling c1susaux has resulted in the optic getting stuck. An approach that seems to work (without getting ITMX stuck)  is to do the following:
    • Save the alignment of the optic, turn off Oplev servo
    • Move the bias sliders on IFO align to (0,0) slowly
    • Turn the watchdog for ITMX off
    • Unplug the cables running from the satellite box to the vacuum feedthrough
    • Power cycle the slow machine. Be aware that when the machine comes back on, the offset sliders are reset to the value in the saved file! So before plugging the cables back in, it would be advisable to set these to (0,0) again, to avoid kicking the optic while plugging the cables back in
    • Plug in the cables, restore alignment and Oplev servos, check that the optic isn't stuck
  4. Y green beat touch up - I tweaked the alignment of the first mirror steering the PSL green (after the beam splitter to divide PSL green for X and Y beats) to maximize the beat amplitude on a fast scope. Doing so increased the beat amplitude on the scope from about 20mVpp to ~35mVpp. A detailed power budget for the green beats is yet to be done

It is unfortunate we have to do this dance each time c1susaux has to be restarted, but I guess it is preferable to repeated unsticking of the optic, which presumably applies considerable shear force on the magnets...


After Wednesday's locking effort, Eric had set the IFO to the PRMI configuration, so that we could collect some training data for the PRC angular feedforward filters and see if the filter has changed since it was last updated. We should have plenty of usable data, so I have restored the arms now.

ELOG V3.1.3-